BASIC ANALOG OF FOURIER SERIES ON A q-QUADRATIC GRID

JOAQUIN BUSTOZ AND SERGEI K. SUSLOV

Abstract. We prove orthogonality relations for some analogs of trigonometric functions on a q-quadratic grid and introduce the corresponding q-Fourier series. We also discuss several other properties of this basic trigonometric system and the q-Fourier series.